Find three sets of values for x and y to make x miles over y minutes = 4 miles over 32 minutes a proportion.
Answer by timvanswearingen(106)

4/32 reduces to 1/8...so one mile per 8 minutes.
2 miles=8*2minutes=16 minutes
3 miles=8*3minutes=24 minutes
and so on